OpenStudy (anonymous):

integral of 1/x^2-4x+5 dx ????

OpenStudy (amistre64):

is that an undernominator?

OpenStudy (amistre64):

\[\frac{1}{x^2-4x+5}=\frac{1}{(x-2)^2+1}\]

OpenStudy (amistre64):

looks to tan-1 up

OpenStudy (anonymous):

how did you get from 1/x^2−4x+5 to 1/(x−2)^2+1 ?

OpenStudy (amistre64):

...in algebra we learn about completing a square

OpenStudy (anonymous):

ohhhhhh! Thank you :)

OpenStudy (amistre64):

yep, when the bottom doesnt factor nicely for a decomp, try to complete the square so you can tan-1 up it